# Equivariant Symmetry-Aware Head Pose Estimation for Fetal MRI

### v1

@@ 1 Introduction
Real-time 6-DoF object pose estimation enables a wide range of critical navigation tasks in robotic manipulation [billard2019], autonomous driving [chiu2021], and image-guided surgery [yaniv2006]. Here, we present a novel method for rapid object pose estimation from 3D images based on an architecture that captures pose equivariance and object symmetries by construction. We are motivated by the problem of motion estimation in fetal MRI, where stacks of 2D diagnostic MRI slices are acquired to assess fetal development and detect pathology (Appendix A). Due to fetal motion, oblique (i.e., misaligned with the target anatomical orientation) slices and spatial coverage gaps in the acquired stack present challenges to radiological assessment and often necessitate re-acquisition of the entire stack (Fig. 1) [glenn2009, malamateniou2013, gholipour2014, xu2023, white2010]. One solution is to estimate the fetal head pose from rapid, low-resolution, low signal-to-noise ratio (SNR) 3D MRI “navigator” volumes inserted into the MRI sequence before every 2D slice. The pose estimated from each navigator volume can be used to adaptively adjust the imaging plane of the next 2D slice in real time (Fig. 1) [hess2011, tisdall2012, gagoski2016]. s Estimating fetal head pose from navigator volumes is particularly challenging. In addition to low resolution and low SNR, these volumes contain an imaging artifact that presents as a dark plane in the position of the previous (high energy) diagnostic 2D slice and may obstruct relevant anatomical features (e.g., eyes) that are crucial for pose disambiguation. Moreover, underdeveloped neuroanatomy coupled with low resolution creates a left-right symmetry, introducing pose ambiguities [zhan2013]. Finally, the lack of available navigator data necessitates training on higher resolution, higher SNR, artifact-free research volumes, thus creating cross-domain generalization challenges.
@@ 2 Related work
Rotational and reflectional object symmetries can lead to pose ambiguities. One class of solutions estimates a range of symmetrically equivalent poses during training, and heuristically selects the most probable pose during inference [pitteri2019, hodan2020, zhao2023]. Yet, these methods suffer from discontinuities in the network output over the object symmetry group [richterklug2021]. Although training with symmetry-invariant losses partially alleviates this problem [xiang2018], these discontinuities are more robustly eliminated by incorporating object symmetry into the pose parametrization itself. Namely, this strategy regresses an output that remains invariant under the object symmetry group, while maintaining continuity over the quotient group of all non-symmetric input transformations [saxena2007, richterklug2021]. To handle head pose ambiguities, E(3)-Pose models the basis direction along the axis of left-right symmetry as a pseudovector [hauser1986], which maintains both invariance under left-right reflections and continuity over SO(3). While pseudovectors have been used to model rotation axes in pose estimation [lee2019], our method is the first to use them to address symmetric ambiguities.
@@ 3 Preliminaries
Let $f\!:\!\mathbb{R}^{3}\!\rightarrow\!\mathbb{R}^{d}$ be a $d$-dimensional tensor field. Consider 3D Euclidean group actions $g\!=\!g_{r}\circ g_{t}$, which comprise rotations/reflections $g_{r}\!\in\!\text{O(3)}$ and translations $g_{t}\!\in\!\mathbb{R}^{3}$. When the coordinate system is transformed by $g$, the tensor $f(x)$ at point $x\!\in\!\mathbb{R}^{3}$ is modified in two ways: it is moved to the location $g^{-1}x$ and the tensor is independently rotated/reflected by $g_{r}$ [weiler2018, geiger2022]. Formally, we can define an operator $\pi_{f}(\cdot)$ that transforms field $f$ for any $g\!=\!g_{r}\circ g_{t}\!\in\!\text{E(3)}$:
@@
$F(\pi_{f}(g)f)=\rho_{h}(g_{r})F(f),\text{ for all }g\!=\!g_{r}\circ g_{t}\in\text{E(3)}.$ (3)
@@ 4.2 Rotation estimation
$\mathcal{L}_{\phi}$ $=\sin{(\theta_{x})}+\beta(\sin{(\theta_{y}/2)}+\sin{(\theta_{z}/2})),$ (9)
@@ 5.4 Results
$\mathcal{L}_{\phi}=\!\sum_{k\in\{x,y,z\}}\sin{(\theta_{k}/2)}$ 8.8 $\mathbf{\pm}$ 3.6 3.0 $\mathbf{\pm}$ 1.3 12.1 $\pm$ 11.1* 7.2 $\pm$ 5.1* 11.1 $\mathbf{\pm}$ 5.3 3.6 $\mathbf{\pm}$ 1.5 18.5 $\pm$ 27.3 9.4 $\pm$ 9.4
@@ Appendix A Automated slice prescription in fetal brain MRI
[…]

### v2

@@ 1 Introduction
Real-time 6-DoF object pose estimation enables a wide range of critical navigation tasks in robotic manipulation [billard2019], autonomous driving [chiu2021], and image-guided surgery [yaniv2006]. Here, we present a novel method for rapid object pose estimation from 3D images based on an architecture that captures pose equivariance and object symmetries by construction. We are motivated by the problem of motion estimation in fetal MRI, where stacks of 2D diagnostic MRI slices are acquired to assess fetal development and detect pathology [glenn2009, gholipour2014] (Appendix A). Due to inter-slice fetal motion, oblique (i.e., misaligned with the target anatomical orientation) slices and spatial coverage gaps in the acquired stack present challenges to radiological assessment and often necessitate re-acquisition of the entire stack (Fig. 1) [malamateniou2013, xu2023, white2010]. One solution is to estimate the fetal head pose from rapid, low-resolution, low signal-to-noise ratio (SNR) 3D MRI “navigator” volumes inserted into the MRI sequence before every 2D slice. The pose estimated from each navigator volume can be used to adaptively adjust the imaging plane of the next 2D slice in real time (Fig. 1) [hess2011, tisdall2012, gagoski2016]. Estimating fetal head pose from navigator volumes is particularly challenging. In addition to low resolution and low SNR, these volumes contain an imaging artifact that presents as a dark plane in the position of the previous (high energy) diagnostic 2D slice and may obstruct relevant anatomical features (e.g., eyes) that are crucial for pose disambiguation. Moreover, underdeveloped neuroanatomy coupled with low resolution creates a left-right symmetry, introducing pose ambiguities [zhan2013]. Finally, the lack of available navigator data necessitates training on higher resolution, higher SNR, artifact-free research volumes, thus creating cross-domain generalization challenges.
@@ 2 Related work
Rotational and reflectional object symmetries can lead to pose ambiguities. One class of solutions trains a network to implicitly learn object symmetries via probabilistic outputs, and heuristically selects the most probable pose during inference [hodan2020, zhao2023]. Yet, these methods suffer from discontinuities in the network output over the object symmetry group [richterklug2021]. Although training with symmetry-invariant losses partially alleviates this problem [xiang2018], these discontinuities are more robustly eliminated by incorporating object symmetry into the pose parametrization itself. Namely, this strategy regresses an output that remains invariant under the object symmetry group, while maintaining continuity over the quotient group of all non-symmetric input transformations [pitteri2019, saxena2007, richterklug2021]. To handle head pose ambiguities, E(3)-Pose models the basis direction along the axis of left-right symmetry as a pseudovector [hauser1986], which maintains both invariance under left-right reflections and continuity over SO(3). While pseudovectors have been used to model rotation axes in pose estimation [lee2019], our method is the first to use them to address symmetric ambiguities.
@@ 3 Preliminaries
Let $f\!:\!\mathbb{R}^{3}\!\rightarrow\!\mathbb{R}^{d}$ be a $d$-dimensional tensor field. Consider 3D Euclidean group actions $g\!=\!g_{t}\circ g_{r}$, which comprise rotations/reflections $g_{r}\!\in\!\text{O(3)}$ and translations $g_{t}\!\in\!\mathbb{R}^{3}$. When the coordinate system is transformed by $g$, the tensor $f(x)$ at point $x\!\in\!\mathbb{R}^{3}$ is modified in two ways: it is moved to the location $g^{-1}x$ and the tensor is independently rotated/reflected by $g_{r}$ [weiler2018, geiger2022]. Formally, we can define an operator $\pi_{f}(\cdot)$ that transforms field $f$ for any $g\!=\!g_{t}\circ g_{r}\!\in\!\text{E(3)}$:
@@
$F(\pi_{f}(g)f)=\rho_{h}(g_{r})F(f),\text{ for all }g\!=\!g_{t}\circ g_{r}\in\text{E(3)}.$ (3)
@@ 4.2 Rotation estimation
$\mathcal{L}_{\phi}$ $=|\sin{(\theta_{x})}|+\beta(|\sin{(\theta_{y}/2)}|+|\sin{(\theta_{z}/2})|),$ (9)
@@ 5.4 Results
$\mathcal{L}_{\phi}=\!\sum_{k\in\{x,y,z\}}|\sin{(\theta_{k}/2)}|$ 8.8 $\mathbf{\pm}$ 3.6 3.0 $\mathbf{\pm}$ 1.3 12.1 $\pm$ 11.1* 7.2 $\pm$ 5.1* 11.1 $\mathbf{\pm}$ 5.3 3.6 $\mathbf{\pm}$ 1.5 18.5 $\pm$ 27.3 9.4 $\pm$ 9.4
@@ Appendix A Automated slice prescription in fetal brain MRI
Radiological assessment of abnormal fetal brain development relies on volumetric fetal brain MRI, which provides greater contrast and resolution than standard ultrasound examinations. Due to sensitivity to fetal motion, current clinical practice does not involve direct acquisition of 3D MRI scans [glenn2009]. Instead, stacks of 2D MRI slices are acquired in the three target anatomical orientations (i.e., sagittal, coronal, and axial). Slices that closely align with the target anatomical orientation are critical for abnormality detection. For instance, non-oblique midline sagittal slices are necessary for detecting agenesis of the corpus callosum [gholipour2014]. Furthermore, inter-slice fetal motion is often correlated with poor spatial coverage. Even small coverage gaps can substantially reduce diagnostic potential. For example, the proportion of the total fetal brain volume constituted by the cerebellum ranges from only 0.03-0.06 in second- and third-trimester fetuses [cai2020]. Adaptive slice prescription based on the fetal head motion thus holds potential to improve the quality of prenatal care.
Slices are T2-weighted and acquired with the half-Fourier acquisition single-shot turbo spin-echo (HASTE) MRI sequence ($1.25\!\times\!1.25$$m$ pixels, 3mm slice thickness, $320\!\times\!320$mm FOV, TR=2.5s, TE=106ms, GRAPPA R=2, Partial Fourier=5/8, $90^{\circ}$ flip angle). Fig. 6 shows examples of diagnostic 2D MRI slices in the target anatomical orientations, as well as oblique slices caused by fetal motion. To automatically prescribe the acquisition plane of each 2D slice according to the current fetal head pose, we rapidly acquire a 3D navigator volume before each slice. Navigator volumes are acquired with the echo-planar imaging (EPI) sequence (4-6mm isotropic voxels, $324\!\times\!324\!\times\!120$mm FOV, TR=29ms, TE=14ms, 5 flip angle). Low-resolution, low-energy navigator volumes are not of diagnostic quality but suffice for head pose estimation.